A book is pushed 3.769 along a horizontal table with a horizontal force of 3N. The opposing force of friction is 0.6N

a.)How much work doesbthe 3N force do on the book?
b.)What isbthe work done on the book by the friction force?
c.)What is the total work done on the book?
1
Expert's answer
2019-11-26T10:35:57-0500

a)


"W_1=F_1d=(3)(3.769)=11.307\ J"

b)


"W_2=-F_2d=-(0.6)(3.769)=-2.261\ J"

c)


"W=W_1+W_2=11.307-2.261=9.046\ N"